OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 9 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase? 96π in3 27π in3 65π in3 105π in3 @zaibali.qasmi

OpenStudy (anonymous):

@SolomonZelman help

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

do you understand?

OpenStudy (kainui):

Imagine the vase without the marbles, can you find the volume then?

OpenStudy (anonymous):

i first got 96

OpenStudy (kainui):

How?

OpenStudy (anonymous):

=pi x 3 x 3 x 12 =108pi next, the volume of one sphere = 4/3 x pi x r^3 (here , d=2 so , r=1) =4/3 x pi x 1 x 1 x 1=4pi/3 so the volume of water will be 108pi-4pi/3 x 9 (9=total no. of marbles) =96pi that is first option

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes..its right.